If an ounce of gold, valued at $610, increases at a rate of 6.6 percent per year, how long will it take to be valued at $1,000? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Number of years
Answer #1
Initial value 610
Amount 1000
rate 6.60%
Life N
1000= 610*(1+6.6%)^N
1000/610= 1.066^N
1.639344262 1.066^N
Log 1.63944 N*log 1.066
0.21467016= N* 0.0277572
N= 0.21467016/0.0277572
N=                  7.73